OpenStudy (anonymous):

Divide the Polynomials (a^2n - a^n - 6) / (a^n+8)

OpenStudy (nikvist):

\[a^{2n}-a^n-6=a^{2n}+8a^n-9a^n-72+66=\]\[=a^n(a^n+8)-9(a^n+8)+66=(\underbrace{a^n-9}_Q)(a^n+8)+\underbrace{66}_R\]